OpenStudy (anonymous):

What is the GCF of the numerator and denominator in the following fraction? 21m^2n/28mn^2 a) 7m^2n^2 b) 7mn c) 21m^3n^3 d) mn

OpenStudy (anonymous):

21m^2 n =3*7*m*m*n 28mn^2 =2*2*7*m*n*n so, GCF = 7*m*n=7mn

OpenStudy (anonymous):

got it?